The following data are available for Sellco for the fiscal year ended on January 31, 2017:

Sales 820 units
Beginning inventory 270 units @ $ 4
Purchases, in chronological order 320 units @ $ 5
400 units @ $ 6
220 units @ $ 8

Below answers were rounded to the nearest dollar.

Cost of Goods Sold Ending Inventory
FIFO 4060 2780
LIFO 5160 1680
Weighted average 4635.37 2205.63

A. Assume that net income using the weighted-average cost flow assumption is $14,300. Calculate net income under FIFO and LIFO. (Round unit cost to 2 decimal places.)

Net Income
FIFO
LIFO
Answer #1

Calculate net income under FIFO and LIFO

Sales revenue under weighted average method = 14300+4635 = $18935

Net income
FIFO 18935-4060 = $14875
LIFO 18935-5160 = $13775